What do you mean???
What is the effect of the anti-whitening filter?
You should know that if you use OUT16 channel, the effect of the unwhite filter is not taken into account.
I made scripts/LSC/LSCoffsets2.py which is the script to zero the dark offset of all the LSC PD. The list of PDs is same as the list in scripts/LSC/LSCoffsets. New script average all outputs of PDs parallelly, so we can zero the offsets much faster.
You can define the averaging time, and you can choose the channel for getting the dark offset from INMON or OUT16. You should know that if you use OUT16 channel, the effect of the unwhite filter is not taken into account.
Example usage (at scripts/LSC):
./LSCoffsets2.py -d 20 --out16
you can find the help by calling this script with option -h or --help
I have modified one of the spare demod boards that was sitting above the electronics bench (the one which was unlabeled - the others say 33MHz, 55MHz and 165MHz) to be the new AS110 demod board. In place of the T1 coil, and the C3 and C6 resistors, I have put the commercial splitter PSCQ-2-120+. In place of U5 (the low pass for the PD input) I have put an SCLF-135+.
In order to figure out how to make the pinout of the PSCQ match up with the available pads from T1, I first pulled the "AS11" board (it's not something that we use, so it would be less of a tragedy if something happened while I had the board pulled). However, while the PCB layout is the same, the splitter for the low frequencies (PSCQ-2-51W) has a different pinout than the one I need for the 110MHz. So, I put AS11 back, and pulled the POP110 board. (After I noted the pinout on POP110, I reinstalled that board. To get it out, I had to unplug the I and Q outs of POP22, but I have also replugged those in).
For my new AS110 demod board, I copied the pin connections on POP110. I have made a little diagram, so you can see what pins went where. The top 2 rectangles are the "before" installation cartoon, and the bottom is the "as installed" cartoon.
The one thing that must be noted is that, because of the pinout of the splitter and the constraints of the board layout, the +0 degrees (I-phase) output of the splitter is connected to the Q channel for the rest of the demod board. This means that the +90 degrees (Q-phase) output of the splitter is connected to the I channel for the rest of the demod board. This is not noted for POP110, but is true for both: The I and Q channels of the 110 MHz demod boards are switched. In practice, we can handle this with our digital phase rotation.
Daytime tomorrow, I will test my new board as Suresh did in elog 4736. Before we get to use AS110, we need (a) some LO juice from the RF distribution box, and (b) a spot to plug the board in, in the LSC rack. Meditating on how those are going to happen are also tasks for daytime tomorrow.
We can put our scripts for the MICH, PRMI, and DRMI into the IFO CONFIGURE screens for now and then it should be easy to get them into the Guardian once Jamie has the bugs worked out.
This screen can also be used to setup and start the dither alignment for each configuration (once we have one working for DRMI / SRM).
Also, now that the notches/bandstop filters for the violin modes have been move from the SUS into the LSC, we should fix the triggering to engage them a few seconds after the boosts.
Friday night locking
Much more stable DRMI lock was achieved, partly thanks to the Friday-night quiet seismic,
and partly because of the improved servo gain and LF boosts
I wanted to confirm the enhancement of the 110MHz signal at the AS port.
As the AS110 PD is placed in the CCD path, there is nothing visible with PRMI.
The Thorlabs PD was moved to the main AS path. Now the AS110 PD is receiving 50% of the power.
With PRMI 110MHz peak was -30dBm (As it was fluctuating, anything more precise number did not make sense)
When the DRMI was locked, the peak was enhanced to 0dBm.
The 2f signal comes from the beat between the sidebands.
Thus the amplitude of the intensity is proportional to the power of the sidebands (assuming the +1 and -1 order sidebands have the same amplitude)
-30dBm -> 0dBm means 31.6 times amplitude of the intensity. Therefore the amplitude transmission of the sidebands is 5.6 times more. (Is this true?)
According to the wiki, the AS port thru-put (i.e. power transmission) for the 55MHz sideband is 0.0026 and 0.43 for PRMI and DRMI respectively.
This corresponds to the amplitude difference of ~13. So we still have only half of the sidebands leaking out from the IFO. This could be attributed
to both the smaller PR gain and SR gain.
Same as the one Jenne used the other day. Later I engaged several additional triggers.
The following is the trigger setting I used
MICH: Delay 2 sec, FM1/FM2/FM3/FM6/FM7
PRCL: Delay 0.5 sec, FM2/FM3/FM6
SRCL: Delay 5 sec, FM1/FM2/FM3/FM6
SRCL FM1 was modified from +3dB to +6dB
Once lock is acquired, it lasts tens of minutes. (see the attached striptool chart.)
Even the lock is lost, it reacquires quickly.
The videos to show the lock acquisition and the in-lock stability are attached below.
The AS port beam is very round. It is not so shaky, but some yaw motion is visible.
The mode at the AS port is defined by the SRM, putting a QPD at the AS port would help to
stabilize the spot.
IFO state upon leaving
I left the 40m with the arms aligned, PRM and SRM slightly misaligned, and LSC setting is for the DRMI locking.
- AS110I/Q for triggering
- PRCL/MICH/SRCL normalization
- We should resurrect the IFO config scripts.
- Remove BS->SRCL actuation coupling
- Handing off to 3f signals (preparation for the full lock)
- Improve ALS stability
- SRM ASC: AS QPD for SRM control
Lock Acquisition Video
UL (REFL) / UR (POP)
LL (AS) / LR (PRM Face)
UL (REFL) / UR (POP)
LL (AS) / LR (PRM Face)
This post describes how to use the Automated Photodetector Frequency Response System.
On the mechanical side, turn on:
-the diode laser (in rack 1Y1)
-the RF Switch (in rack 1Y1)
-the reference PD (under the POY table)
-the AG4395A Network Analyzer
The NA’s RF output should go to the laser’s modulation input, the reference PD’s output should go to the NA’s R input, and the RF Switch Chassis’s output (which is the combination of the two switches’ COM channels using a splitter) should go to the NA’s A input.
Once this is done, navigate into /users/alex.cole and run PDFR.sh. This script collects data for each photodetector under consideration by switching using a python script and communicating with the NA via GPIB. It then sends all the data to RF.m, which fits the functions, plots the latest data against canonical data, and saves the plots to file.
The fitting function, fit.m, also outputs peak frequency to the command line. This function uses PD name data (e.g. ‘REFL33’) to choose an interval with minimal noise to fit.
The main script prompts the user to press enter after each NA sweep to make sure that measurements don’t get interrupted/put out of order by RF switching.
Once you're done, you should turn off the laser, NA, RF Switch, and reference PD.
Sometimes, the NA throws up and doesn’t feel like running a particular sweep. If this happens, it’s a good idea to keep the matlab script from trying to analyze this PD’s data. Do this by opening up RF.m and commenting out the calls to ‘fit’ and ‘canonical’ for that PD.
If fit.m complains about a particular set of data, it is often the case that the N/P ratio (where N is order of approximation and P is number of points in the interval) is too high. You can fix this by reducing N or making the PD’s frequency range (chosen in the fnew_idx line) larger.
Choosing a single PD
If you only want to grab the transfer function for one PD, first look up which switch input it belongs to. This information is contained in /users/alex.cole/switchList. To turn the switch to a particular input, type something like:
python rf.py “ch7”
This command uses TCP/IP to tell the switch to look at channel 7. Switch input numbers range from 1 to 16, though not all of them are in use.
Once the switch is looking at the correct input, you can run a sweep and download the data by typing /opt/rtcds/caltech/c1/scripts/general/netgpibdata/NWAG4395A -s 1000000 -e 500000000 -c 499000000 -f [filestem for output] -d [path of directory for output] -i 192.168.113.108 -g 10 -x 15.
Just to rephrase somewhat:
We were meditating a little bit on what may be the story behind the PRM violin filter situation. We locked the PRMI, and turned on and off the violin filters. We noticed, very bizarrely, that when the violin filters were ON, the servos would oscillate. Weird. Also, probably because the oscillation was causing us to hit the limit we have in the MICH servo, we rung up a 3rd harmonic of one of the violin modes, which was at 1955 Hz.
We took a transfer function of the PRCL servo, saw that the UGF was 300 Hz, and lowered it to ~180 Hz. After later investigations, that high-ish UGF probably wasn't a problem. Anyhow, we then took MICH servo transfer functions, and saw some very weird stuff.
At frequencies where we had violin filter notches, we were seeing peaks in the transfer function, which came close to touching, or crossed the 0dB line! We suspect that this may have something to do with the balancing of the drives to the optics, since we have PRCL driving PRM, but MICH driving BS and PRM. What we did was move the violin filter notches into the LSC model. There were already SUS filter banks in the LSC model (right side of the LSC screen). In preparation for the DRMI, I have put the BS violin notches into the BS, PRM and SRM filter banks, as well as the PRM and SRM filters into all 3 banks. Right now for PRMI, I have the BS and PRM notches (as well as the Vio3 notch) turned on in both BS and PRM. All of the violin-related filters are turned off in the LSC filter bank inside the SUS models. When we did this, the servo oscillations no longer are excited when we turn on the notches, and when we take a new transfer function, there are no longer weird peaks at the notch frequencies. More meditation needs to be done.
Also, for the violin3 filter, Rana noted that at 1955 Hz, after we confirmed that the REFL 55 phase was set correctly (and we're using REFL 55 I&Q for PRMI locking), the I-phase had a response of 0.36, while the Q-phase had a response of 0.20. I should be able to think about these numbers, and decide if the vio3 is for the BS or the PRM violin mode.
the above series of Bode plots shows the MICH Open Loop Gain as the REFL55 phase is adjusted (PURPLE, ORANGE) with the notches in the SUS and then (RED) as the notches are moved to the LSC and made the same for all optics.
In other news, I have the parts that Jamie ordered to make a new 110 demod board, so that's one of my daytime activities now, so we can have both POP110 and AS110.
Meh. 600 counts is too weak. You should fix the electronics so that the maximized green laser transmission gives more like ~10000 counts.
The Y arm green transmission has been measuring in counts all along. I modified the gain in the ALS-TRY filter module to normalise the transmission.
Transmission has been normalised with GTRY = 1 corresponding to 600 counts.
We measured the the openloop transfer function of the PDH green lock of the y-arm.The measurement setup was same as yesterday's measurement.elog 9047
In this measurement, the servo gain was 7 and the source amplitude for the excitation was 1 mV. As you can see in below figure, the measured UGF was 15 kHz and the phase margin was 45 degree.
attatchment1 - OLTF with servo gain of 7
Don't go for a hacky solution. We want to climb a staircase step by step.
Prepare an independent 110MHz demod ports.
To-do: Set up the AS OSA. Also, perhaps temporarily borrow the 110 demod board from POP. We were triggering on POP22 tonight, and that seemed to work okay.
Very nice!! I was wondering, shouldn't the driving matrix be such that MICH pushes on SRM as well?
Hmmm, yes, that's a very good point. I think you're right, and I'll give that a try today.
Wonderful ! I like the video -- the spatial mode looks pretty clean and much cleaner than what I observed in the old days.
The DRMI has been locked!! And at least one time, it was for more than one minute!!
We are not 100% sure yet that it's correctly sideband locked. The test of this was to put a 50% BS in front of the AS camera (so after the beam has gone to AS55), and send the light over to a PDA10CF Thorlabs PD. I locked the Michelson on carrier for the alignment of this diode. Then I strung a cable to the control room, and plugged it into the RF spectrum analyzer. (First, I had turned off the green beat PD power, so there wasn't any RF stuff on the line that I unplugged). It's hard to watch the screen and a tv / dataviewer at the same time, so I've taken a video, so that we can see the nicely locked round DRMI beam on the AS camera, and the spectrum analyzer. My phone is working very hard at uploading the video, but we may have to wait until tomorrow for that. However, I think that we're locked on the 55MHz sideband. (Also, maybe I'm too tired or excited or something, but how do you make the real cameras take video??)
EDIT: Video uploaded. Pause the video at 10 seconds, and you'll see that we've got a strong 110MHz peak!! Hoooray! The TV in the upper right side of the video is AS. You can see as we flash, the peaks go up and down. When there's no resonance, the 110 peak goes away. (Ex., when I'm PRMI locked on the sideband, there isn't a visible peak).
Alignment procedure was as normal: Lock and align the arms. Misalign ETMs. Check that MICH fringes look good (ASS does a nice enough job that I don't actually lock and align the Michelson anymore). Restore the PRM. Lock PRMI. Tweak PRM alignment to maximize POP110I. At this point, Koji and I played a little with the PRMI, but when we finished with that, we restored the SRM, and tweaked its alignment by making nice overlap on the AS camera.
Then, we tried some DRMI settings, started seeing some locks, and played a bit with trying to optmize the settings that we have.
PRCL ASC is on (with loop triggering). MICH gain = -0.8, PRCL gain = +0.05. FM4, FM5 always on, FM2 triggered. Loop and filter module triggering on POP22I. No power normalization. MICH and PRCL locked on REFL55 I&Q, with 1's in the LSC input matrix. PRCL actuating on PRM with +1, MICH actuating on BS with +0.5, PRM with -0.267.
I took transfer functions between REFL55 I&Q and REFL11 I&Q, to determine the relative gains and signs. REFL11I's gain should be -18dB relative to REFL55I, with the opposite sign. We tried PRMI locking with MICH = 1*REFL55Q and PRCL = -0.125*REFL11I for the input matrix. Still no power normalization (we haven't used power norm at all today, so I'll quit writing that).
I took transfer functions between REFL55 I&Q and REFL33 I&Q. REFL33I's gain is -8dB relative to REFL55I, but they have the same sign. We tried locking PRMI with MICH = 1*REFL55Q and PRCL = +0.6*REFL33I. Success.
Next up, some Optickle simulations, to help us go in the right direction for DRMI locking. I checked the signs of the error signals REFL55I (PRM sweep), REFL11I (PRM sweep) and REFL55Q (MICH sweep) in both PRMI and DRMI configurations. For all of these cases, the signs were the same (i.e. no sign flips needed to happen for DRMI locking, relative to PRMI locking). I checked the sensing matrices for DRMI and PRMI for those same signals, and took the ratios of the sensing matrix elements. This gave me the ratio of optical gains for each error signal, in the DRMI case vs. PRMI case, so any servo gain changes should be the inverse of these numbers. These numbers are all DRMI/PRMI: REFL55I PRCL response = 0.76, REFL11I PRCL response = 0.99, REFL55Q MICH response = 18. So, when trying to lock the DRMI, we wanted to keep the gains for PRCL about the same, reduce the servo gain for MICH by a factor of ~20, but keep the same signs for everything.
In doing that, we started seeing some short DRMI locks, so we twiddled some parameters (mostly the elements in the LSC input matrix) a bit. We eventually settled on: PRCL = -0.125*REFL11I, MICH = 0.1*REFL55Q, and SRCL = 1.0 * REFL55I. The output matrix was the same (MICH pushing on BS and PRM, PRCL on PRM), with the addition of a +1 in the SRCL -> SRM element. For all 3 degrees of freedom (PRCL, MICH, SRCL), FMs 4 and 5 were always on. For PRCL, FMs 2,3,6 were triggered to come on after 0.5 seconds of delay. The PRCL FM triggers helped enormously. I tried several other things, including changing the MICH input matrix element up and down in value, changing the SRCL input matrix element up and down in value, and engaging triggering for a few different filters in the MICH and SRCL degrees of freedom. However, none of these made things better, and several made things worse. Most notably, for SRCL, engaging triggering for FMs 2 and 3 kicked the cavities out of lock, which implies that perhaps our gain isn't high enough yet (and thus our UGF isn't very high yet). I changed FM1 of SRCL to be +3dB of gain (from +10dB), and it would live through that coming on (trigger delayed by 1 sec, then ramping up over 1 second), but within a second after the filter finishing coming on, the cavity would fall out of lock (not violently kicked, just not locked anymore).
At this point, we were trying to figure out a way to confirm what kind of lock we had. I checked Optickle again, and we do not expect to see a significant change in POP110I between the PRMI and DRMI cases, so that isn't a useful check. We dreamed of having our AS110 demod board, or the AS OSA set up, but neither of those was going to happen tonight. Instead, Koji suggested hooking up the PD, and looking directly at the output.
To-do: Set up the AS OSA. Also, perhaps temporarily borrow the 110 demod board from POP. We were triggering on POP22 tonight, and that seemed to work okay.
It seems that the PRM violin mode freqs shifted from 625-ish to 640Hz.
The peaks rang up because of the servo.
Once the notch freq was shifted to 640Hz, the violin mode started to decay.
Today we measured OLTF of PDH green lock of x-arm again. In the previous measurement the excitation signal was injected at the PDH servo box output(elog 9044), but in this measurement we changed the injection point to the RFPD mixer output (just before the servo input).
We measured the OLTF with the servo gain of 6.5 and source amplitude of 5 mV for all frequency band. The measured UGF was 11 kHz and the phase margin was 48 degree.
Next that measurement, we tried to measure the power spectrum density of the error signal and feedback signal. But the alignment was not so good, so we aligned the green light injection point. Tomorrow we will continue the alignment and will measure the PSD.
attatchment1 - OLTF of PDH green lock with servo gain of 6.5
Yes, this was not ELOG'd by me, unfortunately. This was the MC tickler which I described to some people in the control room when I turned it on.
As Koji points out, with the MCL path turned off this injects frequency noise and pointing fluctuations into the MC. With the MCL path back on it would have very small effect. After the pumpdown we can turn it back on and have it disabled after lock is acquired. Unfortunately, our LOCKIN modules don't have a ramp available for the excitation and so this will produce some transients (or perhaps we can ezcastep it for now). Eventually, we will modify this CDS part so that we can ramp the sine wave.
I've written a new TICKLE script using the newly found 'cavget' and 'cavput' programs. They are in the standard epics distribution as extension binaries. They allow multichannel read/write as well as ramping, delays, incremental steps, etc. http://www.aps.anl.gov/epics/tech-talk/2012/msg01465.php.
Running from the command line, they seem to work fine, but I've left it OFF for now. I'll switch it into the MC autolocker at some point soon.
One of the reasons that our disk is getting full is due to the scripts_archive directory. A backup script runs on op340m and makes a tar.bz2 file of the scripts directory and puts it in scripts_archive every morning at 6 AM.
On Oct 7, 2011, Koji fixed this script to point at our new scripts directory instead of the old /cvs/cds/caltech/scripts directory. Since then, however, no one has fixed the exclude file to NOT back up the junk that's in that directory. Its a 1.6 GB directory so its full of it.
I've deleted a bunch of junk from the scripts directory: this directory is for scripts, not for your personal home movies or junk data files. Put those in your USER directory. Put temporary data files in /tmp/. I've also added a few more patterns to the exclude file so that less .mpg, .png, .pdf, .dat, etc get stored every day. The new daily .tar.bz2 file wil be ~25 MB instead of 770 MB.
(also fixed the backup script to use 'env' to setup the perl environment and removed the hard-coded path to tar)
Today we measured the openloop transfer function of the PDH green lock of the x-arm.
Edit //manasa// The excitation was given from SR785 source. SR560 was used as the summing node at the PDH servo box output where the loop was broken to measure the OLTF. The SR785 was used to measure the frequency response (CH2/CH1; CH1 A SR560 output and CH2 A PDH servo output) in sweptsine mode.
We measured with two different servo gain. We started with the servo gain of 3 and at that gain the UGF was 1.5 kHz and the phase margin was 50 degree. After that we increase the servo gain to 5.5 and at that gain the UGF was 6.2 kHz and the phase margin was 55 degree. In all the measurement we use the source amplitude of 1.0 mV for all frequencies (from 100 Hz to 100 kHz). We could not increase the gain and also the source amplitude any more because the green was kicked out of lock.
Next work list
1. In the earlier measurements we found the UGF of the PDH green lock of the x-arm as 10 kHz and the phase margin as 45 degree, so we will investigate what has changed from these measurements.elog 4490
2. We will measure the power spectrum of the error signal and the feedback signal.
3. We will calibrate the above signals to compare with ALS out of loop noise.
netgpib was taking forever to transfer data. So the measurements are just photos of the display.
attachment1 - servo gain 3
attachment2 - servo gain 5.5
I have done the swap in the REFL path. First, I swapped the positions of REFL11 and REFL55. Then, I swapped out the 50/50 BS for a 90% reflection BS. (90% goes to REFL55, 10% goes to REFL11). I also changed the aluminum dump that was dumping the old REFL165 path into a razor dump.
Before: REFL11 had 4.0mW, REFL55 had 3.1mW. Now, REFL11 has 0.53mW, and REFL55 has 6.9mW. REFL165 still has around 61mW of light, and REFL33 has 3.3mW (the things that were changed were after 165 and 33 in the REFL path).
Now, the DC value of the REFL PDs are: REFL165 = 10.4V, REFL33 = 110mV, REFL55 = 232mV, REFL11 = 18.6mV.
As I was finishing aligning the beams onto all of the REFL diodes, Manasa asked for the IFO so she and Masayuki could continue their work on the Xarm, so I'll check the signals acquired a little later.
/home/cds is >98% full - below are some of the usage numbers:
controls@rosalba:/users/OLD 0$ du -h --max-depth=1
controls@rosalba:/opt/rtcds/userapps 0$ du -h --max-depth=1
linux1:cds>nice du -h --max-depth=1
du: `./llo/chans/daq/archive': Permission denied
du: `./llo/chans/daq/old': Permission denied
As I always tell everyone: Don't use a 10% reflector which produce ghost beams. Use a 90% reflector.
Hmmm, yes, I forgot (bad me). I'll find a 90% refl BS, and swap the positions of REFL11 and REFL55.
As I always tell everyone: Don't use a 10% reflector which produce ghost beams. Use a 90% reflector.
Shutter moved, no more clipping.
Pick-off mirror 2" replaced by 1" one. Laseroptik HR 532nm, incident angle 30-45 degrees, AR 532 nm
Green REFL PD moved to 4" close to pick-off mirror. Pd being close to pick-off does not separate multiple reflections on it. I'll replace Laseroptic mirror with Al one. It is not easy to find.
Hole cut into side wall for doubler oven cable to exit.
Beam trap for Pd refl is in place. Cabeling is ti·died up.
Laseroptic 1" mirror is replaced by Al 1" mirror. Problem remains the same. This diffraction patter has to be coming from the Faraday.
Atm1, good separation when Pd is far
Atm2, bad separation when Pd is close
According to the wiki, REFL 11 has a transimpedance of 4.08kV/A, and REFL 55 has a transimpedance of 615V/A. This is a ratio of ~6.5 . My optickle simulations from earlier this evening indicate that, at maximum, there is a ~factor of 2 more signal in REFL 11 than REFL 55. This is a factor of order 10-15. Then, REFL 55 has 15dB whitening gain, which is a factor of ~4. So, this explains why we're seeing so much more digital signal on REFL11 than REFL55.
Tomorrow, I need to replace the 50/50 beam splitter that splits the beam between REFL55 and REFL11 (33 and 165 have already had their light picked off at this point). I want to put in a 10% reflector, 90% transmission beamsplitter. Steve, can you please find me one of these, and if we don't have one, order one? This will give us a little more light on 55, and less light on 11, so hopefully we won't be saturating things anymore.
While Jenne was plotting, I locked and aligned the MICH with AS55_Q. Then I aligned the PRM and locked PRMI using REFL55_I/Q with triggering on POP22, but no power normalization.
I used this to set the phase for REFL11 and REFL55 (driving PRM at 111.3 Hz and minimizing the Q response using the DTT Sine Response tool). I flipped the sign on REFL11 by
The REFL11 gain is ~50x larger than REFL55; this is with the 15 dB whitening gain on REFL55 and none for REFL11. What's going on here? The attached PDF shows the two time series with the free swinging PRMI and both phases set to ~ +/- 2 deg. The REFL55 signals have been scaled up by 50x.
So then we went in and looked at the RF signals at the demod boards. To do this we disconnected the RFPD test cables and hooked the RF Mon outputs into the 50 Ohm inputs on a scope. The following PNG images show the scope traces. The REFL11 (yellow) traces are too big!! See how small the REFL55 (green) are. REFL11 is saturating - need to fix.
Here are a bunch of sensing signals. The configuration is always DRMI. Except for the optic noted in the title and the x-axis of any individual plot, other optics are held in their nominal position. DRMI condition is sidebands resonant in PRCL, 55MHz sideband resonant in SRCL. Each plot has an error signal, as well as the 2f signals at POP and AS.
The phases of POP22 and POP110 have been adjusted so that the I signal is maximized when everything is at the nominal positions (sideband resonant for PRMI). The phase of AS110 has been adjusted so that the I signal is maximized when the DRMI is in the nominal position (f2 resonant in SRC). The phases of the 1f1, 1f2, 2f1 and 2f2 REFL signals were all adjusted to have max PRCL signal in the I phase. AS55 was adjusted to have max SRCL signal in the Q phase.
- An Aluminum mirror instead of 2" unknown mirror for the pick-off for the rejected beam from the green faraday isolator (Steve)
=> Replaced. To be reviewed
- Faraday mount replacement. Check what we have for the replacement. (Steve)
- The green REFL PD should be closer to the pick-off mirror. (Steve)
=> Moved. To be reviewed
- A beam dump should be placed for the green REFL PD
- Move the green shutter to the place where the spot is small (Steve)
=> Moved. To be reviewed.
- The pole of the PZT mounting should be replaced with a reasonable one. (Steve with Manasa's supervision)
- Tidying up doubling oven cable. Make a hole on the wall. (Steve)
=> Done. To be reviewed.
- Tidying up the PZT cabling (Steve)
- The optics are dirty. To be drag wiped. (Manasa, Masayuki)
ASX scripts for PZT dither have been fixed appropriately. Script resides in scripts/ASX.
You can run the scripts from the ASX medm screen now.
This disturbance in the MC ASC channels were fixed.
This craziness happened ~10pm last night. Was there any action at the time? >> Sunday-night workers? (RXA: No, Nakano-kun and I left before 9:30 PM)
We found that the signals came from c1ioo. However, restarting, recompiling c1ioo and c1mcs didn't help
to clean up this issue. Just in case we cleaned up the corresponding entries in the ipc file /opt/rtcds/caltech/c1/chans/ipc/C1.ipc
and recomplied c1ioo and c1mcs because these are the channels we touched last week to mitigate the timing out issue of c1rfm.
Incidentally, we fell into a strange mode of the RCG: IOPs could not restart. We ended up running "sudo shutdown -r now"
on each machine (except for c1lsc which was not affected by this issue). This solved the issue.
Even now c1oaf could not be running properly. This is not affecting the IFO operation right now, but we need to look into this issue again
in order to utilize OAF.
When I came in this morning, I noticed that the Mode Cleaner had not been locked for at least the past 8 hours. We moved the MC SUS sliders until the MC SUSPIT and SUSYAW values for each mirror were back to approximately the place they were the last time the MC was nicely locked (~12 hours ago). This got the MC flashing TEM00, so we thought we were doing well.
However, if the servo was enabled, any time the cavity flashed a small-order mode (especially 00), the mirrors would get super kicked. Not good.
We went to investigate, and discovered that the RFPD aux laser was left on again. We turned that off, however that didn't fix the situation.
Manasa suggested checking that the WFS were really, really off. When we looked at the WFS master screen, we noticed that although the WFS servos were off, the MC mirrors' ASC filter banks had non-zero inputs. We checked, and this is not from the MCASS, nor is it from the MC WFS lockins. At this point, I have no idea where these signals are coming from. I have turned off the ASC outputs for all the MC mirrors (which means that we cannot turn on the WFS), and the MC locks fine.
So, we need to know where the ASC signals are coming from. There isn't anything that I can see, from any screen that I can find, that indicates some signals being sent over there. Has anyone done anything lately? I know Koji was working on IPC stuff the other day, but the MC was locking fine over the weekend until yesterday afternoon, so I suspect that's not the culprit.
I have turned off the outputs of the WFS lockins, as part of my turning things off, so if whatever script needs them doesn't enable them, they should be turned back on by hand.
We can not leave cables connected like this. This is a burned toast award.
Masayuki Nakano, a student of Seiji's from ICRR / U Tokyo, is visiting us here at the 40m lab for the next couple months.
He received 40m specific basic safety training this morning.
This is an elog about the activity on Friday night.
- The X arm green beam was aligned with assist of the ASX system.
- M1 PZT alignment was swept while M2 PZT was under the control of ASX.
- Everytime M1 was touched, M2 was restored by manual alignment so that the REFL beam hits the center of the REFL PD.
This way we could recover the lock of TEM00. Once TEM00 is recovered, ASX took care of the alignment of M2
- The error signal used by the cavity dither did not give us a good indication where the optimal alignment is.
- Thus the best alignment of M1 had to be manually scanned. The resulting maximum green transmission was ~0.88
- Once the beam was aligned, the out-of-loop stability of the Xarm was measured.
There has been no indication of the improvement compared to Manasa's measurement taken before our beam alignment.
ETMX sus damping restored
JoeB and JamieR are working somewhat coherently on a set of python libraries to fulfill all of our command line CDS wants. This is being done mostly to satisfy The Guardian and the SkunkTools project.
I did an 'svn up' in /opt/rtcds/userapps (it might finish in ~1000 years) to get the things that they have so far (in particular, Joe's 'pyavg'). There's going to be some issues since the pylib stuff written by Yuta/Kiwamu has never been integrated with anything and is imported as 'epics' in many python scripts. As we move over to the new stuff there will be a lot of broken script functions since the new libraries are also used in that way.
I noticed at LLO (?) that the LSC screen there uses up ~25-30% of the CPU time on a single core for the control room iMac workstations - this seems excessive.
Here is an accounting of CPU usage percentages for some of our screens:
These were measured using the program 'glances' on rosalba. MEDM running with only the sitemap used up 0.9% of a CPU. With the screens running, the fluctuation from sample to sample could be ~ +/- 0.5%. While the LSC screen seems to be the biggest pig, it is only big in comparison to small pigs. Certainly this pig has gotten bigger after getting sent to Louisiana.
Sun Aug 18 15:52:50 2013
Found the FB lights (C1:FEC-NN_FB_NET_STATUS and C1:DAQ-DC0_C1XXX_STATUS) RED for everything on the CDS_FE_STATUS screen.
I used the (! mxstream restart) button ro restart the mxstreams. Everything is green now.
PMC was out of lock- relocked it and the IMC locked itself as did the X & Y arms on IR. X was already green locked.
I took the "aso-laptop" and made it into Ubuntu a couple months ago. Today I added it to the Martian network and then moved it to the X End.
I followed the instructions in (https://wiki-40m.ligo.caltech.edu/Network) and added it to the files in /var/named/chroot/var/named on linux1 and did the "service named restart".
The router already had his MAC address in its list (because Yoichi was illegally using his personal laptop on the Martian). The new laptop's name is 'asia'. This is a legal name according to our computer naming conventions and this Wikipedia page (http://en.wiktionary.org/wiki/Category:Italian_female_given_names). It has been added to the Name Pool on the wiki.
The terminal on the laptop still calls itself 'aso-laptop' so I need some help in fixing that. It successfully connects to 40MARS and displays a MEDM sitemap after sshing in to pianosa.
I use 'ssh -X -C' since I find that compression actually helps when the laptops are so far from the router.
Rana and I connected the PMC_trans output to the BNC connector board on the west end of the PSL table (the channel is labeled). I took a few spectra off of PMC_trans and the SR785 was connected directly to the PMC_trans output for about an hour.
Data will follow.
Yesterday we cleaned up the ASX model and screens to have more straight forward structure of the screen
and the channel names, and to correct mistakes in the model/screens.
The true motivation is that I suspect the excess LF noise of the X arm ALS can be caused by misalignment
and beam jitter coupling to the intensity noise of the beat. I wanted to see how the noise is affected by the alignment.
Currently X-end green is highly misaligned in pitch.
- Any string "XEND" was replaced by "XARM", as many components in the system is not localized at the end table.
- The name like "XARM-ITMX" was changed to "XARM-ITM". This makes easier to create the corresponding model for the other arm.
- There was some inconsistency between the MEDM screens and the ASX model. This was fixed.
- A template StripTool screen was created. It is currently saved in users/koji/template as ASX.stp.
It will be moved to the script directory once it's usefulness is confirmed.
The next step is to go to the end table and manually adjust M2 mirror while M1 is controlled by the ASX.
The test mass dithering provides the error signal for this adjustment but the range of the PZT is not enough
to make the input spot position to be controlled. In the end, we need different kind of matching optics
in order to control the spot position. (But is that what we want? That makes any PZT drift significantly moves the beam.)
Apparently all of the ION pump valves (VIPEE, VIPEV, VIPSV, VIPSE) opened, which vented the main volume up to 62 mTorr. All of the annulus valves (VAVSE, VAVSV, VAVBS, VAVEV, VAVEE) also appeared to be open. One of the roughing pumps was also turned on. Other stuff we didn't notice? Bad.
Several of the suspensions were kicked pretty hard (600+ mV on some sensors) as a result of this quick vent wind. All of the suspensions are damped now, so it doesn't look like we suffered any damage to suspensions.
CLOSE CALL on the vacuum system:
Jamie and I disabled V1, VM2 and VM3 gate valves by disconnecting their 120V solenoid actuator before the swap of the VME crate.
The vacuum controller unexpectedly lost control over the swap as Jamie described it. We were lucky not to do any damage! The ion pumps were cold and clean. We have not used them for years so their outgassing possibly accumulated to reach ~10-50 Torr
I disconnected_ immobilized and labelled the following 6 valves: the 4 large ion pump gate valves and VC1, VC2 of the cryo pump. Note: the valves on the cryo pump stayed closed. It is crucial that a warm cry pump is kept closed!
This will not allow the same thing to happen again and protect the IFO from warm cryo contamination.
The down side of this that the computer can not identify vacuum states any longer.
This vacuum system badly needs an upgrade. I will make a list.
While I was doing the oil change of the roughing pumps I accidentally touched the 24 V adjustment knob on the power supply.
All valve closed to default condition. I realized that the current indicator was red at 0.2A and the voltage fluctuated from 3-13V
Increased current limiter to 0.4A and set voltage to 24V I think this was the reason for the caos of valve switching during the VME swap.
Based on the facts above I reconnected VC1 and VC2 valves. State recognition is working. Ion pumps are turned off and their gate valves are disabled.
We learned that even with closed off gate valves while at atmosphere ion pumps outgass hydrocarbons at 1e-6 Torr level. We have not used them for this reason in the passed 9 rears.
I need help with implementing V1 interlock triggered by Maglev failure signal and-or P2 pressure.
MEDM screen agrees with vacuum rack signs.
After many, many moons of getting to know exactly how frustrating Altium can be, I have completed the PCB layout for my ISS board (final page of ISS_v3.pdf).
Before I get into detail about the PCB, there is one significant schematic change to note: the comparator circuit was changed (with significant help from Koji) so that the voltage reference for boost triggering is established in a more logical way. Instead of the somewhat convoluted topology I had before, now there are only two feedback resistors, R82 and R83. Because their resistances (500k and 50k respectively) are so much larger than the total resistance of the 1k potentiometer (used to establish a tunable threshold voltage), the current flowing through the feedback loop is negligible compared to the 5 mA current flowing through the potentiometer (the pot is rated for 2 W and with 5 mA -> 25 mW dissapation). This allows one to set the threshold voltage for my schmitt trigger, at pin 2 of both the pot and the comparator, entirely with the pot. This trigger also has hysteresis given by the relation deltaV ~ (R83/R82) * (Voh - Vol) where deltaV is the separation between threshold voltages, Voh is the high-level comparator ouput and Vol is the low-level comparator output. Koji simulated this using CircuitLab and I plan to verify the behavior by making a quick prototype circuit.
Now, on to the PCB. The board itself is of a 'standard' LIGO size (11" x 6") has 3 routing layers and 3 internal planes, one for +15 V, one for -15 V and one for GND. In the attached pdf, red is the top routing layer, blue is the bottom layer and brown is the middle routing layer (used for ±5 V exclusively). The grey circles are pads and vias (drilled through) and anything in black is silkscreen overlay. I placed each component and track by hand, attempting to minimize the signal path and following the general rules below,
Sections of the board have been partitioned and labeled with silkscreen overlay to help in both signal pathway recognition as well as eventual troubleshooting.
On the board, I have also included holes so that it can be mounted inside of an enclosure. There is a DCC number printed as well as a 'barcode' (TrueType font: IDAutomationC39S), although they both contain filler asterisks as I haven't published this to the DCC and thus do not have a number.
Beat notes were recovered for both the arms.
I locked the arms to IR using PDH and measured the ALS out of loop noise at the phase tracker output.
The Y arm has the same 300Hz/rtHz rms. The X arm rms noise measures nearly the same as the Y arm in the 5-500Hz region (X arm has improved nearly 10 times after the last whitening filter stage change old elog ).
The noise in the ALSX error signals could be related to the bad alignment and conditions at the X end.